Zeronknight (zeronknight):

Write 0.3% as a ratio in lowest terms. * 3/10 3/1,000 3/100 3/1

sammixboo (sammixboo):

Do you know what 30 percent is equal to?

sammixboo (sammixboo):

Do you know what 30 percent is equal to?

Zeronknight (zeronknight):

No

Zeronknight (zeronknight):

WOndering if it's A

sammixboo (sammixboo):

30% = 3/10, so it isn't, A, but this will give you something to base it off of. Let me give you an example. 60% = 6/10 6% = 6/100 0.6% = 6/1,000 0.06% = 6/10,000 When you divide 3 by 10, you get the decimal 0.3 To find the percent, you move that decimal point over two places to the right, and throw on a percentage sign

sammixboo (sammixboo):

If that makes any sense

Zeronknight (zeronknight):

0.003?

Zeronknight (zeronknight):

i mean it's B, but we didn't exactly move it 2 decimals to the right?

sammixboo (sammixboo):

0.3% = 0.003, yes. Since I told you how to convert a fraction to a percent, you do the exact opposite when turning a percent to a fraction.

Zeronknight (zeronknight):

Move it two decimals to the left, actually making it 0.3%

sammixboo (sammixboo):

Therefore, you move it two places to the left

Zeronknight (zeronknight):

aka 3/1000

Zeronknight (zeronknight):

Very helpful. thank you :)

sammixboo (sammixboo):

LOL sorry, I got myself mixed up with all of these 0's

sammixboo (sammixboo):

But you understand, correct?

Zeronknight (zeronknight):

Yes, If I wanted to convert this to a percent then fraction, like 0.05% IT would become 5/1000?

sammixboo (sammixboo):

It would become 10,000

sammixboo (sammixboo):

5/10,000*

Zeronknight (zeronknight):

Oh XD

sammixboo (sammixboo):

;) Yeah, I have it memorized by a chart, so it's harder for me to explain it with words

Zeronknight (zeronknight):

Thank you :P